### v1

@@ 3 Results
fixed effect on age, fixed effect on education, IID by county Female 0.31
fixed effect on age, RW1 by education, IID by county Female 0.31
fixed effect on age, RW1 by education, BYM2 by county Female 0.31
@@
fixed effect on age, fixed effect on education, IID by county Male 0.33
fixed effect on age, RW1 by education, IID by county Male 0.33
fixed effect on age, RW1 by education, BYM2 by county Male 0.33
fixed effect on age, BYM2 by education Male 0.34
Results comparing modelled estimates aggregated from the county- to the state-level with CDC baseline estimates are visualized in Figure 1 and reported in Table LABEL:tab:state_propestimates in the Appendix. State-level aggregates for all four models captured the direct mean survey estimates produced from CTIS survey weights, but did not capture the CDC estimate for the proportion of individuals in California who had received the first-dose COVID-19 vaccination for June 30, 2021.
Figure 1: Results of the number of first-dose COVID-19 vaccinations modelled for both sexes at the county level and aggregated to the state level produced using classic MRP methods (IID by county), and spatial MRP methods (BYM2 models). Models were able to capture the direct survey estimate of first-dose COVID-19 vaccinations at the state level produced by the CTIS survey weights, but did not capture the CDC estimate for June 30, 2021.
Results for the county-level estimates produced in this analysis are illustrated in Figure 2 and reported in Table LABEL:tab:county_propestimates in the Appendix. We recall the weights provided in the CTIS are known to be inappropriate for producing vaccination estimates below the state level. We use the county-level CDC estimates of first-dose vaccination for June 30, 2021 as the closest approximation of the truth with which to compare the validity of our modelled estimates. MRP estimates would be considered accurate if they captured CDC state-level estimates (and even more desirably, county-level estimates). In this case, spatial and classic MRP performed similarly and did not capture CDC estimates, tending to over-smooth estimates downward.
@@ 4 Discussion
MRP is a quick, computationally inexpensive, and relatively simple-to-implement indirect estimation method. Its ease and flexibility has made it increasingly popular. New advancements in statistical computation now allow us to implement both classic and spatial MRP pipelines for extremely complex Bayesian hierarchical models using INLA. We can now produce and visualize these estimates in minutes. From an implementation standpoint, this is extremely exciting.
@@
As with any statistical method, MRP has limitations. In this analysis, we were unable to capture the “true” county-level estimates of first-dose vaccination provided by the CDC, though all four models contained state-level estimates produced by the direct mean survey weights. Unfortunately, MRP is a method that is only effective when its model accurately reflects the outcome of interest in the target population (Kennedy and Gelman, 2021). It will produce estimates for virtually all demographic strata across each geographic unit; it is up to us to validate these. In modelling situations, we make decisions based on expertise provided in the appropriate literature and also according to variables present in the available data. While MRP is able to make predictions for all observed and unobserved areas represented in the survey data used to train the model, the model will be unable to effectively capture the outcome of interest if there is inadequate population and contextual data available with which to predict and poststratify.
@@
Here, we followed the classic MRP approach of only including covariates present in the data; these were limited because they had to match categorization definitions across data sources (i.e., same age ranges used for age bins, etc.). As previously noted, it has been suggested that COVID-19 vaccine hesitancy in California (and the US, more broadly) had been driven by lower likelihood of vaccination among members of groups historically marginalized according to race/ethnicity (e.g., Nguyen et al., 2022; Khubchandani and Macias, 2021; Willis et al., 2021), but by early 2021 COVID-19 vaccine trends in California were significantly driven by and positively associated with greater educational attainment (Thomas et al., 2021). As the CDC did not release a cross-tabulation of county vaccination counts by sex, age, race/ethnicity, and education, we could not predict the probability of these strata and reliably verify them. Modelling choices in MRP are necessarily based on strong assumptions with respect to demographic structures that underlie the outcome of interest in the target population. Excluding key structuring demographics from the model due to a lack of availability or granularity in the poststratification data suffuses the model with unintended but strong, perhaps even incorrect, assumptions. This was the case with this evaluation: first-dose vaccination estimates are likely to be strongly structured according to sex, age, educational attainment, but this analysis provides evidence that the demographics-specific outcome data represented in this survey sample do not capture important subpopulation vaccination trends. These models were also limited because CDC estimates were only available for very few age categories 18-24 years, 26-64 years, 65+ years).
@@
Naturally, the question remains: how could we improve these models, and MRP estimation broadly? Kastellec et al. (2015) suggest expanding the poststratification table by using auxiliary survey data including area-level ecological/population characteristics with the aim of reducing the differences between the survey sample and target population. In our case, including race/ethnicity population proportions may improve vaccination estimates, but it remains that the validation data do not contain these details. As Kennedy and Gelman (2021) remind us, we must know our populations and models to produce generalizable results beyond our samples. To this end, another area of improvement is to try to improve the survey sample selection process. As the CTIS was administered through the Facebook platform, user activity drove sample selection, where more frequently active users were more likely to be selected. Therefore, the sample selection process contained an unavoidable temporal process, where in a daily random sample, more active users will of course be oversampled. In this case, the CTIS significantly oversampled individuals in an over-aggregated age category: aged 25-64 years. The problems inherent to this aggregation decision are difficult to overcome with MRP. It would also be beneficial to be provided better data with respect to survey weights. Companies like Facebook are producing large-scale datasets with relatively principled sampling schemes, but they do not provide the core marginals to accurately adjust the survey data they share. We encourage these companies to share their user-representative metrics so we can benchmark estimates to them.
@@
Clearly, there are important limitations to using MRP in the current health and demographics data landscape. Indirect estimation methods are used when direct estimation methods are unavailable, such as in emergency response and planning situations. This is includes public health agencies’ needs with respect to tracking effects and aspects of COVID-19 response during the first years of the pandemic. Indirect estimation methods therefore have the potential to meet a critical need in times of crisis. Better demographics must be recorded and released to researchers to improve local- and state-level response and policy decisions. We must not only work to balance individual privacy concerns with data granularity, but also align demographic strata and geographic data choices in survey products with generalizable population data used in national census products, and also make them available to researchers.
4.1 Limitations
The CTIS itself has a few limitations and errors. The Delphi Group describe these on the CTIS website (Delphi Group, 2024). They acknowledge the survey weights are inappropriate for estimation below the state level. Additionally, between June 15 and July 4, 2021, a random 1% of the sample was sent the instrument for the previous wave (10) rather than that which was current at the time (wave 11). The Delphi Group claim to have not detected trend discontinuities as a result of the error, despite differences in the two instruments. Additionally, between June 17 and 24, 2021, users using an Android phone to access the Facebook application through which the survey was administered were unable to open the app due to a bug in the in-app web browser. The Delphi Group reported that they had recovered 89.5% of the initial decrease in total response volume as of June 26, 2021.

### v2

@@ 3 Results
fixed effect on age, RW(1) by education, BYM2 by county Female 0.313
fixed effect on age, RW(1) by education, IID by county Female 0.313
fixed effect on age, fixed effect on education, IID by county Female 0.314
@@
fixed effect on age, fixed effect on education, IID by county Male 0.333
fixed effect on age, RW(1) by education, BYM2 by county Male 0.333
fixed effect on age, RW(1) by education, IID by county Male 0.333
fixed effect on age, BYM2 by education Male 0.343
Results comparing modelled estimates aggregated from the county- to the state-level with CDC baseline estimates are visualized in Figure 1 and reported in Table LABEL:tab:state_propestimates in the Appendix. State-level aggregates for all four models captured the mean survey estimates produced from CTIS survey weights, but did not capture the CDC estimate for the proportion of individuals in California who had received the first-dose COVID-19 vaccination for June 30, 2021.
Figure 1: Results of the number of first-dose COVID-19 vaccinations modelled for both sexes at the county level and aggregated to the state level produced using classic MRP methods (IID by county), and spatial MRP methods (BYM2 models). Models were able to capture the mean survey estimate of first-dose COVID-19 vaccinations at the state level produced by the CTIS survey weights, but did not capture the CDC estimate for June 30, 2021.
Figure 2: Results of the number of first-dose COVID-19 vaccinations modelled by sex and age aggregated to the state level produced using classic MRP methods (IID by county), and spatial MRP methods (BYM2 models).
When we consider a finer grain visualization of estimates by age and sex (Figure 2), we see all models over-estimate vaccination for both sexes and all age groups. We also see an underlying issue with the CDC estimate for males and females aged 65 years and older is greater than the total population. We expect this is a result of either undocumented older individuals residing in California, or else because non-residents of California were vaccinated in the state. These results imply the CDC baseline estimate at the state level in Figure 1 is higher than the true proportion of Californians vaccinated on June 30, 2021.
Results for the county-level estimates produced in this analysis are illustrated in Figure 3 and reported in Table LABEL:tab:county_propestimates in the Appendix. We recall the weights provided in the CTIS are known to be inappropriate for producing vaccination estimates below the state level. We use the county-level CDC estimates of first-dose vaccination for June 30, 2021 as the closest approximation of the truth with which to compare the validity of our modelled estimates. MRP estimates would be considered accurate if they captured CDC state-level estimates (and even more desirably, county-level estimates). In this case, spatial and classic MRP performed similarly and did not capture CDC estimates consistently, tending instead to overestimate vaccination prevalence; random effects models smooth with data they are provided, but are unable to handle this level of bias in the data.
@@
3.1 Limitations
The CTIS itself has a few limitations and errors. The Delphi Group describe these on the CTIS website (Delphi Group, 2024). They acknowledge the survey weights are inappropriate for estimation below the state level. Additionally, between June 15 and July 4, 2021, a random 1% of the sample was sent the instrument for the previous wave (10) rather than that which was current at the time (wave 11). The Delphi Group claim to have not detected trend discontinuities as a result of the error, despite differences in the two instruments. Additionally, between June 17 and 24, 2021, users using an Android phone to access the Facebook application through which the survey was administered were unable to open the app due to a bug in the in-app web browser. The Delphi Group reported that they had recovered 89.5% of the initial decrease in total response volume as of June 26, 2021. Another limitation is that while the CDC reported counts of first-dose vaccination are the best available baseline available for comparison between survey and model estimates and the “true” prevalence of first-dose vaccination, these data are imperfect. As we demonstrated in Figure 2, the counts for both males and females ages 65 years and older are greater than the total population in these sex and age groups reported in the ACS 5-year estimates.
@@ 4 Discussion
MRP is a quick, computationally inexpensive, and relatively simple-to-implement estimation method. Its ease and flexibility has made it increasingly popular. New advancements in statistical computation now allow us to implement both classic and spatial MRP pipelines for extremely complex Bayesian hierarchical models using INLA. We can now produce and visualize these estimates in minutes. From an implementation standpoint, this is extremely exciting.
@@
As with any statistical method, MRP has limitations. In this analysis, we were unable to capture the “true” county-level estimates of first-dose vaccination provided by the CDC for the vast majority of counties, though all four models contained state-level estimates produced by the mean survey weights. Unfortunately, MRP is a method that is only effective when its model accurately reflects the outcome of interest in the target population (Kennedy and Gelman, 2021). It will produce estimates for virtually all demographic strata across each geographic unit; it is up to us to validate these. In modelling situations, we make decisions based on expertise provided in the appropriate literature and also according to variables present in the available data. While MRP is able to make predictions for all observed and unobserved areas represented in the survey data used to train the model, the model will be unable to effectively capture the outcome of interest if there is inadequate population and contextual data available with which to predict and poststratify.
@@
Here, we followed the classic MRP approach of only including covariates present in the data; these were limited because they had to match categorization definitions across data sources (i.e., same age ranges used for age bins, etc.). As previously noted, it has been suggested that COVID-19 vaccine hesitancy in California (and the US, more broadly) had been driven by lower likelihood of vaccination among members of groups historically marginalized according to race/ethnicity (e.g., Nguyen et al., 2022; Khubchandani and Macias, 2021; Willis et al., 2021), but by early 2021 COVID-19 vaccine trends in California were significantly driven by and positively associated with greater educational attainment (Thomas et al., 2021). As the CDC did not release a cross-tabulation of county vaccination counts by sex, age, race/ethnicity, and education, we could not predict the probability of these strata and reliably verify them. Modelling choices in MRP are necessarily based on strong assumptions with respect to demographic structures that underlie the outcome of interest in the target population. Excluding key structuring demographics from the model due to a lack of availability or granularity in the poststratification data suffuses the model with unintended but strong, perhaps even incorrect, assumptions. This was the case with this evaluation: first-dose vaccination estimates are likely to be strongly structured according to sex, age, educational attainment, but this analysis provides evidence that the demographics-specific outcome data represented in this survey sample do not consistently capture important subpopulation vaccination trends. These models were also limited because CDC estimates were only available for very few age categories (18-24 years, 26-64 years, and 65+ years).
@@
Naturally, the question remains: how could we improve these models, and MRP estimation broadly? Kastellec et al. (2015) suggest expanding the poststratification table by using auxiliary survey data including area-level ecological/population characteristics with the aim of reducing the differences between the survey sample and target population. In our case, including race/ethnicity population proportions may improve vaccination estimates, but it remains that the validation data do not contain these details. As Kennedy and Gelman (2021) remind us, we must know our populations and models to produce generalizable results beyond our samples. To this end, another area of improvement is to try to ameliorate the survey sample selection process. As the CTIS was administered through the Facebook platform, user activity drove sample selection, where more frequently active users were more likely to be selected. Therefore, the sample selection process contained an unavoidable temporal process, where in a daily random sample more active users will, of course, be oversampled. In this case, the CTIS significantly oversampled individuals in an over-aggregated age category: aged 25-64 years. The problems inherent to this aggregation decision are difficult to overcome with MRP. It would also be beneficial to be provided better data with respect to survey weights. Companies like Facebook are producing large-scale datasets with relatively principled sampling schemes, but they do not provide the core marginals to accurately adjust the survey data they share. We encourage these companies to share their user-representative metrics so we can benchmark estimates to them.
@@
Clearly, there are important limitations to using MRP in the current health and demographics data landscape. Estimation methods are used when directly observed data are unavailable, such as in emergency response and planning situations. This includes public health agencies’ needs with respect to tracking effects and aspects of COVID-19 response during the first years of the pandemic. Estimation methods therefore have the potential to meet a critical need in times of crisis. Better demographics must be recorded and released to researchers to improve local- and state-level response and policy decisions. We must not only work to balance individual privacy concerns with data granularity, but align demographic strata and geographic data choices in survey products with generalizable population data used in national census products, and also make them available to researchers.